Eric's Easter Eggs manufactures large chocolate eggs. The length of these eggs follows a normal distribution with mean = 13 inches and standard deviation = 1.8 inches.

If ONE egg is randomly selected, what is the probability the egg will be longer than 10.8 inches?

Must show ALL work
Answer #1

Solution :

Given ,

mean = = 13

standard deviation = = 1.8

P(x > 10.8) = 1 - P(x<10.8 )

= 1 - P[(x -) / < (10.8-13) /1.8 ]

= 1 - P(z < -1.22)

Using z table

= 1 - 0.1112

= 0.8888

probability= 0.8888